Cathy mixes 5 pounds of dried cranberries that sell for $4 per pound with some dried peaches that sell for $12 per pound. If she wants the mix to be worth $10 per pound, then how many pounds of peaches should she use?
Answer by mananth(16946)

--------price ---------------- quantity
cranberries 4 ---------------- 5 pounds
peaches 12 ---------------- x pounds
Mixture 10 ---------------- 5 + x pounds

4*5+12x =10(5+x)
20+12x= 50+10x
12x+(-10)x=50-(-20)
2 x = 30
/ 2
x = 15
15 pounds of peaches